How many students are offering both biology and chemistry,if only 70 students are offering neither biology or chemistry.


1
Expert's answer
2021-12-23T12:05:53-0500

QUESTION:

Out of 400 students in a final year of a senior secondary, 300 are offering Biology and 190 are offering Chemistry.


How many students are offering both Biology and Chemistry if 70 students are offering neither Biology nor Chemistry.


SOLUTION:

Let "n" "=" the number of students offering both Biology and Chemistry.

Biology student only"=(300-n)"

Chemistry student only"=(190-n)"

Neither Biology nor Chemistry"=70"


"\therefore n+(300-n)+(190-n)+70=400"


"560-n=400\\n=160"